VIBRONIC SPECTROSCOPY OF JET-COOLED 2-FLUORO-m-XYLYL AND 2-CHLORO-m-XYLYL RAD-ICALS GENERATED BY CORONA DISCHARGE

\underline{YOUNG WOOK YOON}, CHANG SOON HUH, SANG KUK LEE, Department of Chemistry, Pusan National University, Pusan 609-735, Korea.

By means of a technique of corona excited supersonic expansion (CESE) combined with a pinhole-type glass nozzle developed in this laboratory, we generated vibronically excited but jet-cooled 2-fluoro-m-xylyl and 2-chloro-m-xylyl radicals from precursors 2-fluoro-m-xylene 2-chloro-m-xylene with a large amount of carrier gas He, respectively. The well-resolved vibronic emission spectra of the jet-cooled radical species were recorded with a long-path monochromator in the visible region. From the analysis of the spectrum, we determine an accurate electronic energy of the D_{1} → D_{0} transition and the frequencies of vibrational modes in the ground electronic state by comparison with those of ab initio calculations and the known spectroscopic data of precursors for the first time. The red-shift of the origin band of each species from parental benzyl radical agrees well with those predicted from synergic effect of each substituent into benzene ring.